A market researcher at a major company classified households by car ownership. The relative frequencies of households for each category of ownership are shown in Table 1.


Table 1. The relative frequencies of households


Number of cars House hold 0 1 2 3 4 5


Relative Frequency 0.1 0.3 0.4 0.12 0.06 0.02



Calculate the mean value and standard deviation of the random variable and interpret the result.



1
Expert's answer
2022-06-15T16:53:39-0400
"mean=\mu=E(X)=0.1(0)+ 0.3(1)+ 0.4(2)"

"+ 0.12(3)+ 0.06(4)+ 0.02(5)=1.8"

"E(X^2)=0.1(0)^2+ 0.3(1)^2+ 0.4(2)^2"

"+ 0.12(3)^2+ 0.06(4)^2+ 0.02(5)^2=4.44"

"Var(X)=\sigma^2=E(X^2)-(E(X))^2"

"=4.44-(1.8)^2=1.2"

"\sigma=\sqrt{\sigma^2}=\sqrt{1.2}\approx1.0954"

A large standard deviation (or variance) means that the distribution is spread out, with some chance of observing values at some distance from the mean.